A population consists of four numbers 2, 3, 4, 5. Consider all possible distinct

samples of size two with replacement. Find (i) the population mean (ii) the

population standard deviation (iii) the sampling distribution of means (iv)

standard deviation of the sampling distribution of means


1
Expert's answer
2021-07-18T09:14:49-0400

(i)


"mean=\mu=\dfrac{2+3+4+5}{4}=3.5"

(ii)

"Variance=\sigma^2"

"=\dfrac{(2-3.5)^2+(3-3.5)^2+(4-3.5)^2+(5-3.5)^2}{4}"

"=1.25"

"\sigma=\sqrt{\sigma^2}=\sqrt{1.25}=\dfrac{\sqrt{5}}{2}"

(iii) We have population values "2, 3, 4, 5," population size "N=4," and sample size "n=2." Thus, the number of possible samples which can be drawn with replacement is



"N^n=4^2=16"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 No & Sample & Mean \\ \hline
 1 & (2, 2) & 2 \\
 \hdashline
 2 & (2,3) & 2.5 \\
 \hdashline
3 & (2,4) & 3\\
 \hdashline
4 & (2,5) & 3.5 \\
 \hdashline
5 & (3,2) & 2.5 \\
 \hdashline
6 & (3, 3) & 3\\
 \hdashline
7 & (3,4) & 3.5 \\
 \hdashline
8 & (3, 5) & 4 \\
 \hdashline
9 & (4,2) & 3 \\
 \hdashline
10 & (4, 3) & 3.5 \\
 \hdashline
11 & (4,4) & 4 \\
 \hdashline
12 & (4, 5) & 4.5 \\
 \hdashline
13 & (5,2) & 3.5 \\
 \hdashline
14 & (5, 3) & 4 \\
 \hdashline
15 & (5,4) & 4.5 \\
 \hdashline
16 & (5, 5) & 5\\
 \hdashline
\end{array}"

The sampling distribution of the sample mean "\bar{x}" and its mean and standard deviation are:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:}
 \bar{x} & f & f(\bar{x})& \bar{x}f(\bar{x})& \bar{x}^2f(\bar{x}) \\ \hline
 2 & 1 & 1/16 & 2/16 & 4/16 \\
 \hdashline
 2.5 & 2 & 2/16 & 5/16 & 12.5/16 \\
 \hdashline
 3 & 3 & 3/16 & 9/16 &27/16 \\
 \hdashline
 3.5 & 4 & 4/16 & 14/16 & 49/16 \\
 \hdashline
 4 & 3 & 3/16 & 12/16 & 48/16 \\
 \hdashline
 4.5 & 2 & 2/16 & 9/16 & 40.5/16 \\
 \hdashline
 5 & 1 & 1/16 & 5/16 & 25/16 \\
 \hdashline
 Total &1 6 & 1 & 3.5 & 103/8\\
 \hdashline
\end{array}"

"E(\bar{X})=\sum\bar{x}f(\bar{x})=3.5"

"Var(\bar{X})=\sum\bar{x}^2f(\bar{x})-(\sum\bar{x}f(\bar{x}))^2"

"=\dfrac{103}{8}-\dfrac{49}{4}=\dfrac{5}{8}"

"E(\bar{X})=3.5=\mu"

(iv)


"Var(\bar{X})=\sum\bar{x}^2f(\bar{x})-(\sum\bar{x}f(\bar{x}))^2"

"=\dfrac{103}{8}-\dfrac{49}{4}=\dfrac{5}{8}"

"Var(\bar{X})=\sigma_{\bar{x}}^2=\dfrac{5}{8}=\dfrac{\sigma^2}{n}=\dfrac{5}{4(2)}"

"\sigma_{\bar{x}}=\dfrac{\sqrt{10}}{4}=\dfrac{\sigma}{\sqrt{n}}"